# On the Generalization Error of Differentially Private Algorithms Via Typicality

### v1

@@ Abstract
We study the generalization error of stochastic learning algorithms from an information-theoretic perspective, with a particular emphasis on deriving sharper bounds for differentially private algorithms. It is well known that the generalization error of stochastic learning algorithms can be bounded in terms of mutual information and maximal leakage, yielding in-expectation and high-probability guarantees, respectively. In this work, we further upper bound mutual information and maximal leakage by explicit, easily computable formulas, using typicality-based arguments and exploiting the stability properties of private algorithms. In the first part of the paper, we strictly improve the mutual-information bounds by Rodríguez-Gálvez et al. (IEEE Trans. Inf. Theory, 2021). In the second part, we derive new upper bounds on the maximal leakage of learning algorithms. In both cases, the resulting bounds on information measures translate directly into generalization error guarantees.
@@ I Introduction
A randomized statistical learning algorithm $P_{W|S}$ is defined as a probabilistic mapping from a training dataset $S:=\left(Z_{1},\ldots,Z_{N}\right)\in\mathcal{Z}^{N}$ consisting of $N$ independent and identically distributed (i.i.d.) data points from distribution $P_{Z}$ (which is unknown) on the space $\mathcal{Z}$ to a hypothesis $W$ from the hypothesis space $\mathcal{W}$. We assume $\mathcal{Z}$ is finite and $\mathcal{W}$ is discrete, like [1]. The generalization performance of this learning algorithm is measured by the difference between the expected loss $\mathbf{E}_{P_{Z}}\left[\ell(W,Z)\right]$ and the empirical loss $\frac{1}{N}\sum^{N}_{i=1}\ell(W,Z_{i})$, where $\ell(\cdot,\cdot)$ is the loss function. We assume $\ell(\cdot,\cdot)$ is $\sigma$-sub-Gaussian in this paper, i.e., for all $\lambda\in\mathbb{R}$ and all $w\in\mathcal{W}$,
@@
One of the key challenges in statistical learning theory is to guarantee that a learning algorithm can “generalize well”, i.e., the generalization error defined as
$\mathrm{gen}(W,S):=\mathbf{E}_{P_{Z}}\left[\ell(W,Z)\right]-\frac{1}{N}\sum^{N}_{i=1}\ell(W,Z_{i}),$ (1)
is bounded and hence how much the algorithm overfits to the data is under control.
A vast array of studies have been proposed for this objective. To guarantee that the trained algorithm can generalize well, the theory of uniform convergence [2] states that the output has to be “sufficiently simple”. Following this idea, bounds in terms of the VC-dimension and the Rademacher complexity have been proposed [3, 4]. However, they make no reference to the algorithm. To make a specific algorithm generalize, various types of methods have been studied, e.g., based on compression schemes [5] or uniform stability [6]. Motivated by controlling the bias in data analysis using the mutual information between the empirical risks of the hypotheses and the algorithm output [7], it was shown in [8] that the expectation of the generalization error can be bounded by the mutual information $I(S;W)$ between the training dataset $S$ and the output hypothesis $W$:
$\mathbf{E}\left[\big|\mathrm{gen}(W,S)\big|\right]\leq\sqrt{\frac{2\sigma^{2}}{N}I(S;W)}.$ (2)
This provides the intuition that “an algorithm that leaks little input information generalizes well” [9, 10], which partially coincides with the idea of private algorithms. There exist various definitions of privacy, including total variation distance [11], differential privacy [12, 13], and max-information [14]. See Section II-B for a detailed discussion.
Maximal leakage $\mathcal{L}(S\rightarrow W)$ [15] of an algorithm $P_{W|S}$, as another information leakage measure, has been proposed and discussed recently [16, 17, 18, 19]. In [18], it has been shown that the generalization error can also be bounded by $\mathcal{L}(S\rightarrow W)$: given $\eta\in(0,1)$,
@@
One major meaningful difference between (3) and (2) is that (3) is a high-probability (a.k.a. single-draw [20, 21]) bound that admits an exponentially decaying (in the size of the training dataset $N$) bound on the probability of having a large generalization error, which is generally untrue for mutual-information-based bounds that only guarantee a linearly decaying bound [10].
For generalization error bounds in terms of other information measures, see [22, 23, 24, 18, 21, 25, 26]. The connection between privacy measures and generalization has been widely discussed in [14, 27, 28, 18, 1, 29]. Note that privacy is tightly related to compression [30, 31], and compressibility can be used to bound the generalization error as well [5, 23, 24].
In [1], an approach based on typicality [32] has been used to bound the generalization error of private algorithms, and their upper bounds are in more explicit forms, since mutual information is sometimes hard to calculate.
In this paper, our contributions are two-fold. In the first part of this paper (Section III), we utilize the stability property of differentially private algorithms to further strengthen the mutual information bounds in [1], and show that our results strictly improve the existing bounds. In the second part of this paper (Section IV), we propose novel upper bounds on the maximal leakage, which directly translate to upper bounds on mutual information as well [15].
Notations
We use calligraphic letters (e.g., $\mathcal{W}$), capital letters (e.g., $W$) and lower-case letters (e.g., $w$) and to denote sets, random variables and instances, respectively. We assume all random variables are from finite alphabets and logarithm and entropy are to the base $e$ unless otherwise stated. For a statement $S$, $\mathds{1}_{\{S\}}$ is its indicator, i.e., $\mathds{1}_{\{S\}}$ is $1$ if $S$ holds, and 0 otherwise.
@@ II Preliminaries
We use $S:=\left(Z_{1},\ldots,Z_{N}\right)$ to denote a training dataset, which is a collection of $N$ instances $Z_{i}\in\mathcal{Z}$ i.i.d. from $P_{Z}$. A stochastic learning algorithm $\mathcal{A}:\mathcal{Z}^{N}\rightarrow\mathcal{W}$ can be characterized by the conditional distribution $P_{W|S}$, i.e., $S$ is a random variable with distribution $P_{S}=P_{Z}^{\otimes N}$ over $\mathcal{Z}^{N}$, where $|\mathcal{Z}|$ is finite. We consider algorithms that are permutation-invariant: $\mathcal{A}$ operates on a set instead of an ordered sequence; the hypothesis generated by $\mathcal{A}$ with dataset $S$ is equal in distribution to the hypothesis generated by $\mathcal{A}$ with a permuted dataset $\mathcal{P}(S)$, i.e., $\mathcal{A}(S)\stackrel{{\scriptstyle d}}{{=}}\mathcal{A}(\mathcal{P}(S))$ for all permutations $\mathcal{P}$. This is a common assumption in the generalization analysis of learning theory; see also [6, 33].
In this paper, we upper-bound the generalization error (1) via either an in-expectation bound of the form (2) or a high-probability bound of the form (3). There are other types of generalization error bounds, e.g., PAC-Bayesian bounds [21, 34, 35], which may also be of interest in the future. We consider $\sigma$-sub-Gaussian loss functions in this paper; note that if the loss function $\ell$ is bounded on $[a,b]$, then it is (ba)/2(b-a)/2-sub-Gaussian.
@@ II-A Method of Types
Definition 1 ([32]).
The type of a sequence $Z^{N}\in\mathcal{Z}^{N}$ is the relative proportion of occurrences of each symbol of $\mathcal{Z}$, i.e., $T_{Z^{N}}(t)=\mathsf{N}(t\,|\,Z^{N})/N$ for all $t\in\mathcal{Z}$, where $\mathsf{N}(t\,|\,Z^{N})$ is the number of times the symbol $t$ occurs in the sequence $Z^{N}$. The set of all possible types of sequences of elements from $\mathcal{Z}$ of length $N$ is denoted by $\mathcal{T}_{\mathcal{Z},N}$.
The type and its variants were used to prove achievabilities of network information theory problems, either asymptotic [36] or nonasymptotic [37, 38]. For the latter, also see other methods [39, 40, 41]. Even though the number of items in $\mathcal{Z}^{N}$ scales exponentially in $N$, it is known that the total number of types only scales polynomially: [32] gives $|\mathcal{T}_{\mathcal{Z},N}|\leq(N+1)^{|\mathcal{Z}|}$, which was marginally improved by [1]:
@@
Moreover, the property of the type that will be utilized in this paper is that it is equal to that of any permutation of the same sequence, i.e., $T_{Z^{N}}=T_{\mathcal{P}(Z^{N})}$ for all permutation $\mathcal{P}$, and hence $T_{Z^{N}}$ can uniquely define a dataset $S=\{Z_{1},\ldots,Z_{N}\}$. Hence [1] can define the type of the dataset $S$ as
$T_{S}\triangleq T_{\tilde{Z}^{N}}$
where $\tilde{Z}^{N}$ is a sequence generated by all the $N$ instances of $S$, and therefore the distance between two datasets $S,S^{\prime}$ can be defined via their types $T_{S},T_{S^{\prime}}$.
@@
where $\mathsf{N}(a|S)$ denotes the number of times the symbol $a$ occurs in the dataset $S$.
@@ II-B Differential Privacy
Differential privacy [12, 13], as one of the most celebrated privatization schemes, has been widely studied in the past years (also see its variants [42, 43, 44]):
@@
Given an algorithm $\mathcal{A}$ which induces the conditional distribution $P_{W|S}$, we say it satisfies $(\varepsilon,\delta)$-DP if for any neighboring (differ in a single data point) $(s,s^{\prime})$ and $\mathcal{Z}\subseteq\mathcal{W}$, it holds that for all measurable $W$,
$\mathbf{P}\big(W\in\mathcal{Z}|S=s\big)\leq e^{\varepsilon}\,\mathbf{P}\big(W\in\mathcal{Z}|S=s^{\prime}\big)+\delta.$
@@
One useful property is group privacy [45, Theorem 2.2]: for an $\varepsilon$-DP algorithm $\mathcal{A}$, if two (fixed) datasets $S$ and $S^{\prime}$ are at distance $k$, then the KL divergence of their output distributions is bounded by
@@
Moreover, for an $\varepsilon$-DP mechanism, in nats, we have [46]
$I(S;W)\leq\min\{\varepsilon,\varepsilon^{2}\}.$
Discussions on maximal leakage can be found in Section IV.
We are now ready to present our main results.
@@ III Mutual Information Bounds
In [1], upper bounds on the generalization error of discrete algorithms have been proposed by upper bounding $I(W;S)$, or more precisely upper bounding the relative entropy $D_{\mathrm{KL}}(P_{W|S}\,\|\,Q_{W})$, since by the golden formula [47, Section 4.1],
$I(W;S)\leq\mathbf{E}_{s\sim P_{S}}\left[D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\right],$ (5)
where the equality is achieved if and only if $Q_{W}=P_{W}$, and the expression $D_{\mathrm{KL}}(P_{W|S}\,\|\,Q_{W})$ also appears in other generalization error characterizations, e.g., [21]. By (2), these upper bounds directly translate to generalization error bounds.
Then based on the typicality result in Claim 1, [1] has the following results:
Proposition 1 ([1]).
@@
    $D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\leq(|\mathcal{Z}|-1)\log(1+e\varepsilon N).$
Combining with the in-expectation generalization error bound (2) and the golden formula (5), from (6) one can derive
$\big|\mathbf{E}_{P_{WS}}\left[\mathrm{gen}(W;S)\right]\big|\leq\sqrt{\frac{2\sigma^{2}}{N}\left(|\mathcal{Z}|-1\right)\log(N+1)}.$
In this paper, we first show a slightly improved version of Claim 1, since the proof of Claim 1 (see [1, Appendix G.1]) used a the very loose step $1+N/i\leq 1+N$ for every $i$.
@@
Our Claim 2 is strictly better than Claim 1 by a constant factor <1<1, and this constant factor $(k+1)/2^{k}$ is actually an optimal factor $c$ such that $\binom{N+k}{k}\leq c(N+1)^{k}$ holds for all $N\geq 1$. Recall $k=|\mathcal{Z}|-1$, it can be dramatically tighter as $|\mathcal{Z}|$ grows. However, we note that the advantages of our theorems do not mostly come from here, we are just using Claim 2 to make our bounds as tight as possible. In the bounds presented later, we will show them utilizing both claims separately.
@@
Let $P$ and $Q$ be two probability distributions such that $P\ll Q$. Let $Q$ be a finite mixture of probability distributions such that $Q=\sum_{b}\omega_{b}Q_{b}$, where $\sum_{b}\omega_{b}=1$, and $P\ll Q_{b}$ for all $b$. Then, the following inequalities hold:
@@
If $Q_{b}=P_{W|S=s_{b}}$ are output distributions of an $\varepsilon$-DP algorithm and $\mathsf{d}(\cdot,\cdot)$ is the dataset distance, then group privacy implies for all measurable $E$,
@@
Hence the lemma applies with $\alpha_{b,i}=e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$ for any $b$.
@@
The bound by Lemma 3 is always at least as good as (8) since $\sum^{M}_{b=1}\omega_{b}\alpha_{b,i}\geq\omega_{i}\alpha_{i,i}=\omega_{i}$, and is strictly better whenever $\sum_{b\neq i}\omega_{b}\alpha_{b,i}>0$.
@@
We first define what do we mean by representative sets, which is from the construction presented in [1, Appendix C]. The idea is to utilize the property that a private algorithm produces statistically similar outputs for neighboring datasets. This will also be utilized in our proof.
Note all types $T_{s}$ lie inside the unit hypercube $0,1]^{|\mathcal{Z}|-1}$, and can be uniquely identified by the preceding $(|\mathcal{Z}|-1)$ dimensions. Rather than discussing types, we study the vector of counts $\mathsf{N}_{s}$ by $\mathsf{N}_{s}(a)=\mathsf{N}(a|s)$ for all $a\in\mathcal{Z}$, and hence $\mathsf{N}_{s}$ lies in a $0,N]^{|\mathcal{Z}|-1}$ hypercube. For the interval $[0,N]$, each dimension can be split to $t$ parts where integer $t\in[1:N]$, hence resulting in a $0,N]^{|\mathcal{Z}|-1}$ hypercube covering of $t^{|\mathcal{Z}|-1}$ hypercube smaller hypercubes, and the side of each small hypercube has length $l=N/t$ and $l^{\prime}=\lfloor l\rfloor+1$ atoms.
The representative sets is constructed by:
@@
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$, with an induced distribution $P_{W|S}$ where $W\in\mathcal{W}$ is a hypothesis obtained with an algorithm $\mathcal{A}$. For all dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}$ on $\mathcal{W}$ such that for every $t\in\{1,2,\dots,N\}$,
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}\bigr)$
$\leq(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+(|\mathcal{Z}|-1)\log t-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon N}\bigr).$ (9)
@@
For the second term, it vanishes when $t=1$, and when $t\geq 2$ we know $M\geq 2$, picking $j\neq i$ gives $\sum_{b}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\geq 1+e^{-\varepsilon N}$. Hence it is at most $\log M-\log(1+e^{-\varepsilon N})\leq(m-1)\log t-\log(1+e^{-\varepsilon N})$. The proof is complete by combining these arguments, and we can minimize over $t$ as well. ∎
@@
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$. Fix $|\mathcal{Z}|\geq 2$ and $N\geq 1$. For every $t\in\{1,2,\dots,N\}$ and every dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}^{(t)}$ such that
@@
$\qquad-\log((|\mathcal{Z}|-1)!)-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr).$ (10)
@@
Furthermore, similar to the idea of [1, Proposition 4] by employing the strong typicality, we can further improve our bounds and derive the following result.

### v2

@@ Abstract
We study the generalization error of stochastic learning algorithms from an information-theoretic perspective, with particular emphasis on deriving sharper bounds for differentially private algorithms. It is well known that the generalization error of stochastic learning algorithms can be bounded in terms of mutual information and maximal leakage, yielding average and high-probability guarantees, respectively. In this work, we further upper bound mutual information and maximal leakage by explicit, easily computable formulas, using typicality-based arguments and exploiting the stability properties of private algorithms. In the first part of the paper, we strictly improve the mutual-information bounds by Rodríguez-Gálvez et al. (IEEE Trans. Inf. Theory, 2021). In the second part, we derive new upper bounds on maximal leakage of learning algorithms. In both cases, the resulting bounds on the information measures translate directly into generalization error guarantees.
@@ I Introduction
A randomized statistical learning algorithm $P_{W|S}$ can be defined as a probabilistic mapping from a training dataset $S:=\allowbreak\left(Z_{1},\ldots,Z_{N}\right)\in\mathcal{Z}^{N}$ consisting of $N$ independent and identically distributed (i.i.d.) data points from distribution $P_{Z}$ (which is unknown) on the space $\mathcal{Z}$ to a hypothesis $W$ from the hypothesis space $\mathcal{W}$. Similar to [1], we assume $\mathcal{Z}$ is finite and $\mathcal{W}$ is discrete. The generalization performance of this algorithm is measured by the difference between the expected loss $\mathbf{E}_{P_{Z}}\left[\ell(W,Z)\right]$ and the empirical loss $\frac{1}{N}\sum^{N}_{i=1}\ell(W,Z_{i})$, where $\ell(\cdot,\cdot)$ is the loss function. We assume $\ell(\cdot,\cdot)$ is $\sigma$-sub-Gaussian [^1] A random variable (r.v.) bounded on $[a,b]$ is $\frac{b-a}{2}$-sub-Gaussian. If $(X_{i})_{i}$ are independent $\sigma$-sub-Gaussian r.v.s, $\frac{1}{n}\sum^{n}_{i=1}X_{i}$ is $\frac{\sigma}{\sqrt{n}}$-sub-Gaussian. in this paper, i.e., for all $\lambda\in\mathbb{R}$ and all $w\in\mathcal{W}$,
@@
One of the central challenges in statistical learning theory is to guarantee that a learning algorithm $\mathcal{A}$ can “generalize well”, i.e., the generalization error defined as
$\mathrm{gen}(W,S):=\mathbf{E}_{P_{Z}}\left[\ell(W,Z)\right]-\frac{1}{N}\sum^{N}_{i=1}\ell(W,Z_{i})$ (1)
is bounded; and hence how much $\mathcal{A}$ overfits is under control.
A vast array of studies have been proposed for this objective. To guarantee that the trained algorithm can generalize well, the theory of uniform convergence [2] states that the output has to be “sufficiently simple”. Similar to this idea, bounds in terms of the VC-dimension and the Rademacher complexity have been discussed [3, 4]. However, they make no reference to the algorithm. To make a specific algorithm generalize, various types of methods have been studied, e.g., based on compression schemes [5] or uniform stability [6]. Motivated by controlling the bias in data analysis using the mutual information between the algorithm input and output, it was shown by [7, 8] that the expectation of the generalization error can be bounded by the mutual information $I(S;W)$ between the training dataset $S$ and the output hypothesis $W$:
$\mathbf{E}\left[\left|\mathrm{gen}(W,S)\right|\right]\leq\sqrt{\frac{2\sigma^{2}}{N}I(S;W)}.$ (2)
This provides the intuition that “an algorithm that leaks little information about the dataset generalizes well” [9, 10], which partially coincides with the property of private algorithms. There exist various notions of privacy, e.g., total variation distance [11], max-information [12] and differential privacy [13, 14]. See Section II-B for a detailed discussion.
Maximal leakage $\mathcal{L}(S\rightarrow W)$ [15] of an algorithm $P_{W|S}$, as an information leakage measure, has been discussed recently [16, 17, 18, 19]. It has been shown by [18] that the generalization error can be bounded by $\mathcal{L}(S\rightarrow W)$: given $\eta\in(0,1)$,
@@
One major meaningful difference between (2) and (3) is that (3) is a high-probability (a.k.a. single-draw [20, 21]) bound; that is, it admits an exponentially decaying (in the size of the dataset $N$) bound on the probability of having a large generalization error, which is generally untrue for mutual-information-based bounds that only guarantee a linearly decaying bound [10]. Moreover, (3) depends on the samples through its support only and hence independent from the distribution, this can be useful in analysis of additive-noise settings [22].
Generalization error bounds in terms of other information measures can be found in [23, 24, 25, 18, 21, 26, 27]. The connection between privacy measures and generalization has been discussed in [12, 28, 29, 18, 22]. Privacy is also tightly related to compression [30, 31], and compressibility can be used to bound the generalization error as well [5, 24, 25]. In [1], an approach based on typicality [32] has been used to bound the generalization error of private algorithms, and their bounds are in more explicit, easily computable forms, since mutual information can sometimes be hard to calculate.
The contributions of the current paper are two-fold. In the first part (Section III), we utilize the stability property of differentially private algorithms to further strictly improve the mutual information bounds in [1]. In the second part (Section IV), we derive novel upper bounds on maximal leakage, which directly translate to upper bounds on mutual information [15]. We hence provide explicit generalization bounds via (2) and (3).
Notation
We use calligraphic letters (e.g., $\mathcal{W}$), capital letters (e.g., $W$) and lower-case letters (e.g., $w$) to denote sets, random variables and instances, respectively. We assume all random variables are from finite alphabets and logarithm and entropy are to the base $e$ unless otherwise stated. For a statement $S$, $\mathds{1}_{\{S\}}$ is its indicator, i.e., $\mathds{1}_{\{S\}}$ is $1$ if $S$ holds, and 0 otherwise. For distributions $P$ and $Q$, we use $P\ll Q$ to denote that $P$ is absolutely continuous with respect to $Q$.
@@ II Preliminaries
We use $S:=\left(Z_{1},\ldots,Z_{N}\right)$ to denote a training dataset, which is a collection of $N$ instances $Z_{i}\in\mathcal{Z}$ i.i.d. from $P_{Z}$ (which is unknown). A stochastic learning algorithm $\mathcal{A}:\mathcal{Z}^{N}\rightarrow\mathcal{W}$ can be characterized by the conditional distribution $P_{W|S}$, i.e., $S$ is a random variable with distribution $P_{S}=P_{Z}^{\otimes N}$ over $\mathcal{Z}^{N}$, where $|\mathcal{Z}|$ is finite. We consider algorithms that are permutation-invariant: $\mathcal{A}$ operates on a set instead of an ordered sequence; the hypothesis generated by $\mathcal{A}$ with dataset $S$ is equal in distribution to the hypothesis generated by $\mathcal{A}$ with a permuted dataset $\mathcal{P}(S)$, i.e., $\mathcal{A}(S)\stackrel{{\scriptstyle d}}{{=}}\mathcal{A}(\mathcal{P}(S))$ for all permutations $\mathcal{P}$. This is a common assumption in the generalization analysis of learning theory [6, 33].
In this paper, we upper-bound the generalization error (1) via either an average bound of the form (2) or a high-probability bound of the form (3). There exist other types of generalization bounds, e.g., PAC-Bayesian bounds [21, 34, 35], which may also be of interest in the future. We consider $\sigma$-sub-Gaussian loss functions in this paper.
@@ II-A Method of Types
Definition 1.
The type of a sequence $Z^{N}\in\mathcal{Z}^{N}$ is the relative proportion of occurrences of each symbol of $\mathcal{Z}$, i.e., $T_{Z^{N}}(t)=\mathsf{N}(t\,|\,Z^{N})/N$ for all $t\in\mathcal{Z}$, where $\mathsf{N}(t\,|\,Z^{N})$ denotes the number of times symbol $t$ occurs in the sequence $Z^{N}$. The set of all possible types of sequences of elements from $\mathcal{Z}$ of length $N$ is denoted by $\mathcal{T}_{\mathcal{Z},N}$.
The type and its variants were used to prove achievability results in network information theory problems, including both asymptotic [36] and nonasymptotic [37, 38] settings. For the latter case, also see other methods [39, 40, 41]. Even though the number of items in the space scales exponentially in $N$, it is known that the total number of types only scales polynomially: [32] gives $|\mathcal{T}_{\mathcal{Z},N}|\leq(N+1)^{|\mathcal{Z}|}$, which was improved by [1]:
@@
Moreover, the property of the type that will be utilized in this paper is that it is equal to that of any permutation of the same sequence, i.e., $T_{Z^{N}}=T_{\mathcal{P}(Z^{N})}$ for all permutations $\mathcal{P}$. Hence, $T_{Z^{N}}$ can uniquely define a dataset $S=\{Z_{1},\ldots,Z_{N}\}$. We then define the distance between two datasets $S$ and $S^{\prime}$ via their types $T_{S}$ and $T_{S^{\prime}}$, respectively.
@@
where $\mathsf{N}(a|S)$ denotes the number of times sample $a$ occurs in the dataset $S$.
@@ II-B Differential Privacy
Differential privacy [13, 14], as one of the most celebrated privatization schemes, has been widely studied in the past years (also see its variants [42, 43, 44]). It is defined as follows:
@@
Given an algorithm $\mathcal{A}$ which induces the conditional distribution $P_{W|S}$, we say it satisfies $(\varepsilon,\delta)$-DP if for any neighboring (differ in a single data point) $(s,s^{\prime})$ and $\mathcal{V}\subseteq\mathcal{W}$, it holds that for all measurable $W$,
$\mathbf{P}\big(W\in\mathcal{V}|S=s\big)\leq e^{\varepsilon}\,\mathbf{P}\big(W\in\mathcal{V}|S=s^{\prime}\big)+\delta.$
@@
One useful property of DP is group privacy [45, Theorem 2.2]: for an $\varepsilon$-DP algorithm $\mathcal{A}$, if two (fixed) datasets $S$ and $S^{\prime}$ are at distance $k$, then the KL divergence of their output distributions is bounded by
@@
Maximal leakage, another privacy measure on which this paper focuses, will be discussed in Section IV.
@@ III Mutual Information Bounds
In this section, we present novel generalization bounds that can be efficiently computed by bounding mutual information.
This idea has been explored by [1], which provides generalization bounds of discrete, private algorithms by bounding the relative entropy $D_{\mathrm{KL}}(P_{W|S}\,\|\,Q_{W})$, since by the golden formula [46, Section 4.1] we know
$I(S;W)\leq\mathbf{E}_{s\sim P_{S}}\left[D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\right],$ (5)
and the equality is achieved if and only if $Q_{W}=P_{W}$. The expression $D_{\mathrm{KL}}(P_{W|S}\,\|\,Q_{W})$ also appears in other generalization error characterizations, e.g., [21]. These upper bounds on relative entropy can be directly translate to generalization error bounds by (2). We first review a result in [1] by Claim 1:
Proposition 1.
@@
    $D_{\mathrm{KL}}(P_{W|S=s}\,\|\,Q_{W})\leq(|\mathcal{Z}|-1)\log(1+e\varepsilon N).$ (7)
Note that in the regime $\epsilon>1$, (6) is tighter than (7).
Combining with the average generalization error bound (2) and the golden formula (5), from (6) one can prove
$\mathbf{E}\left[\left|\mathrm{gen}(W,S)\right|\right]\leq\sqrt{\frac{2\sigma^{2}}{N}\left(|\mathcal{Z}|-1\right)\log(N+1)}$
and a sharper bound
$\mathbf{E}\left[\left|\mathrm{gen}(W,S)\right|\right]\leq\sqrt{\frac{2\sigma^{2}}{N}(|\mathcal{Z}|-1)\log(1+e\varepsilon N).}$
if $\mathcal{A}$ is also $\varepsilon$-DP with $\varepsilon\leq 1$.
In this paper, we first provide a slightly improved version of Claim 1, since the proof of Claim 1 (see [1, Appendix G.1]) used a the very loose step $1+N/i\leq 1+N$ for every $i$.
@@
Claim 2 is strictly better than Claim 1 by a constant factor <1<1. Take $k=|\mathcal{Z}|-1$, this constant factor $(k+1)/2^{k}$ is actually an optimal factor $c$ such that $\binom{N+k}{k}\leq c(N+1)^{k}$ holds for all $N\geq 1$. it can be dramatically tighter as $|\mathcal{Z}|$ grows. However, we note that the advantages of our theorems do not mostly come from here, we are just using Claim 2 to make our bounds as tight as possible.
@@
Let $P$ and $Q$ be two probability distributions such that $P\ll Q$. Let $Q$ be a finite mixture of probability distributions such that $Q=\sum_{b}\omega_{b}Q_{b}$, where $\sum_{b}\omega_{b}=1$, and $P\ll Q_{b}$ for all $b$. Then we have
@@
If $Q_{b}=P_{W|S=s_{b}}$ are output distributions of an $\varepsilon$-DP algorithm and $\mathsf{d}(\cdot,\cdot)$ is the dataset distance, then group privacy implies that for all measurable $E$,
@@
Hence Lemma 3 applies with $\alpha_{b,i}=e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}$ for any $b$.
@@
The bound by Lemma 3 is always at least as tight as the one in (9) since $\sum^{M}_{b=1}\omega_{b}\alpha_{b,i}\geq\omega_{i}\alpha_{i,i}=\omega_{i}$, and is strictly tighter whenever $\sum_{b\neq i}\omega_{b}\alpha_{b,i}>0$.
@@
We first define the concept of representative sets, introduced in [1, Appendix C], which will also be used in our proof. The idea is to utilize the property that a private algorithm produces statistically similar outputs for neighboring datasets.
Note that all types $T_{s}$ lie inside the unit hypercube $0,1]^{|\mathcal{Z}|-1}$ as the type of a sequence can be uniquely identified by the first $(|\mathcal{Z}|-1)$ dimensions. Rather than discussing types, we study the vector of counts $\mathsf{N}_{s}$ by $\mathsf{N}_{s}(a)=\mathsf{N}(a|s)$ for all $a\in\mathcal{Z}$, and hence $\mathsf{N}_{s}$ lies in a $0,N]^{|\mathcal{Z}|-1}$ hypercube. For the interval $[0,N]$, each dimension can be split into $t$ parts, where $t\in[1:N]$, resulting in a $0,N]^{|\mathcal{Z}|-1}$ hypercube covering of $t^{|\mathcal{Z}|-1}$ smaller hypercubes, and the side of each small hypercube has length $l=N/t$ and $l^{\prime}=\lfloor l\rfloor+1$ atoms.
We obtain the representative set by taking one center atom (one dataset) from each small hypercube in $t^{|\mathcal{Z}|-1}$ cover, by:
@@
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$, with an induced distribution $P_{W|S}$ where $W\in\mathcal{W}$ is a hypothesis obtained by an algorithm $\mathcal{A}$. For all datasets $s\in\mathcal{S}$, there exists a distribution $Q_{W}$ on $\mathcal{W}$ such that
$D_{\mathrm{KL}}\bigl(P_{W|S=s}\,\|\,Q_{W}\bigr)\leq\min_{t\in\{1,2,\dots,N\}}\Big\{(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+$
$\qquad\qquad\quad(|\mathcal{Z}|-1)\log t-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon N}\bigr)\Big\}.$ (10)
@@
For the second term, it vanishes when $t=1$, and when $t\geq 2$ we know $M\geq 2$, picking $j\neq i$ gives $\sum_{b}e^{-\varepsilon\mathsf{d}(s_{b},s_{i})}\geq 1+e^{-\varepsilon N}$. Hence it is at most $\log M-\log(1+e^{-\varepsilon N})\leq(m-1)\log t-\log(1+e^{-\varepsilon N})$. The proof is completed by combining these arguments and minimizing over $t$ as well. ∎
@@
Let $\mathcal{A}$ be an $\varepsilon$-DP algorithm under the dataset distance $\mathsf{d}(\cdot,\cdot)$. Fix $|\mathcal{Z}|\geq 2$ and $N\geq 1$. For every dataset $s\in\mathcal{S}$, there exists a distribution $Q_{W}^{(t)}$ such that
@@
$\leq\min_{t\in\{1,2,\dots,N\}}\Big\{(|\mathcal{Z}|-1)\frac{\varepsilon N}{t}+(|\mathcal{Z}|-1)\log\left(t+\frac{|\mathcal{Z}|-2}{2}\right)$
$\qquad-\log((|\mathcal{Z}|-1)!)-\mathds{1}_{\{t\geq 2\}}\log\bigl(1+e^{-\varepsilon\Delta_{t}}\bigr)\Big\}.$ (11)
@@
Furthermore, similar to the idea of [1, Proposition 4] by employing the strong typicality, we can also improve our bounds and derive the following result.